A cross is performed between red and white flowers. All F1 are pink and are self-crossed. 250 progeny are observed, all of which were pink except for 60 red and 60 white. During investigating if flower color follows Mendel’s 1st Law, you calculate a Chi2 value of ___

a. 159.600

b. 23.778

c. 17.330

d. 0.400

e. 212.933

f. 107.320

g. None of the above
Answer #1

the F1 progeny of the red plants and white plants are all pink, so the alleles shows incomplete dominance

let the alleles be R- red, W- white

RR * WW

RW

RW * RW

R W
R RR ( red) RW ( pink)
W RW (pink) WW ( white)

the expected ratio is red: pink: white=1:2:1

the observed ratio is

red: pink:white=60:130:60=1:2.17:1

phenotype Expected (E) observed (O) (E-O)^2/E
red 1 1 0
pink 2 2.17 0.014
white 1 1 0

chi-square value= sum of (E-O)^2/E=0+0.014+0=0.014

so the answer is g) none of the above.